OpenStudy (calculusxy):

Hector waters his lawn if it does rain at least 1 1/2 inches of rain each week. It has rained 3/8 inch this week. Which inequality represents the number of inches, r, Hector's lawn still needs so that he won't have to water it? (I will write the options down in the reply box)

OpenStudy (calculusxy):

\[r \ge 1 \frac{ 1 }{ 7 } \] \[r > 1 \frac{ 7 }{ 8 }\] \[r \ge 1 \frac{ 1 }{ 8 }\] \[r > 1 \frac{ 1 }{ 8 }\]

OpenStudy (calculusxy):

@mathmale

OpenStudy (calculusxy):

@agent0smith

OpenStudy (agent0smith):

find 1 1/2 - 3/8